# Local Dimension Spectrum for Dominated Irreducible Planar Self-Affine Sets

### v1

@@ 6 Multifractal Formalism
In this section we upgrade the symbolic multifractal formalism results of Section 4 to give results in terms of the Hausdorff dimension of the sets
@@
This section contains a number of small propositions, some of which are quite similar to each other, and the proofs of Theorems 1.2 and 1.3 are given after them.
@@
Let $X$ be a dominated planar self-affine set satisfying the strong separation condition, $\mu\in\mathcal{M}(\Sigma)$ be a fully supported quasi-Bernoulli measure, $10<q<$ 1, and $s\geqslant 0$. Then
@@
In general, the result follows from [13, Theorem 6.2(a)] and [19, Proposition 2.5(iv)]. Our version follows directly from [8, Lemmas 5.1 and 5.3] and the comments after the lemmas. The restriction to $10<q<$ 1 is since, in the language of [8, Lemma 5.1], the inequality between $\tau(\mu,q)$ and $\tau(q)$ changes sign at $q=1$. ∎
The arguments of the following proposition are an adjustment of those in [18, Theorem 4.1].
@@
[…]
@@
is a Vitaly covering of $X_{s}$. By the Vitaly covering theorem (see e.g. [11, Theorem 1.10]), the above family contains a countable sub-family $\mathcal{V}$ of pairwise disjoint balls satisfying
@@
Since the elements of $\mathcal{V}$ are pairwise disjoint, a simple volume argument shows that, under the assumption 1, #$\#\mathcal{V}_{\mathtt{j}}\leqslant c$, and under the assumption 2, #$\#\mathcal{V}_{\mathtt{j}}\leqslant c\frac{\alpha_{1}(\mathtt{j})}{\alpha_{2}(\mathtt{j})}$ for some constant $c>0$ not depending on $\mathtt{j}$. Notice that, as $q>0$, the definition of $\mathcal{V}$ yields
@@
By (6.3), the assumption 1, the fact that #$\#\mathcal{V}_{\mathtt{j}}\leqslant c$, and (6.1), we have
@@
for all $\mathtt{i}\in\Sigma$, $V\in X_{F}$, and $n\in\mathbb{N}$. Therefore, by (6.3), (6.4), the fact that #$\#\mathcal{V}_{\mathtt{j}}\leqslant c\frac{\alpha_{1}(\mathtt{j})}{\alpha_{2}(\mathtt{j})}$, and (6.1), we similarly get
@@
The following proposition is a variant of [8, Proposition 5.5], in which a random version of the same statement was proved without assuming irreducibility of the self-affine set, but under the additional assumption that each $\|A_{i}\|<\frac{1}{2}$.
@@
Let $X$ be a dominated irreducible planar self-affine set satisfying the strong separation condition, $\mu\in\mathcal{M}(\Sigma)$ be a fully supported quasi-Bernoulli measure, and $10<q<$ 1. If $\tau(q)/(q-1)$ and $q\tau^{\prime}(q)-\tau(q)$ are contained in $(0,1)$, then
Proof.
Let $\nu\in\mathcal{M}_{\sigma}(\Sigma)$ be the equilibrium state for $\psi^{q,\tau(q)/(q-1)}$. Since $\nu$ is quasi-Bernoulli, (5.1) and Theorem 4.2 imply that
$\operatorname{\underline{dim}_{H}}(\pi_{*}\nu)=\operatorname{dim_{L}}(\nu)=q\tau^{\prime}(q)-\tau(q).$ (6.5)
By Theorem 5.2, there is $0\leqslant t\leqslant 2$ such that
$\operatorname{dim_{loc}}(\pi_{*}\mu,\pi(\mathtt{i}))=t$
for $\nu$-almost all $\mathtt{i}\in\Sigma$ and hence, $\pi_{*}\nu(X_{t})=1$. Since $\nu$ is ergodic and $\mu$ and $\alpha_{1}$ are sub-multiplicative, Lemma 2.1 shows that
$\lambda_{1}(\nu)$ $=\lim_{n\to\infty}\frac{1}{n}\log\alpha_{1}(\mathtt{i}|_{n}),$
$h(\mu,\nu)$ $=-\lim_{n\to\infty}\frac{1}{n}\log\mu([\mathtt{i}|_{n}])$
for $\nu$-almost all $\mathtt{i}\in\Sigma$. Therefore, Proposition 4.1 implies
$\begin{split}t&=\lim_{n\to\infty}\frac{\log\pi_{*}\mu(B(\pi(\mathtt{i}),\operatorname{diam}(X)\alpha_{1}(\mathtt{i}|_{n})))}{\log\alpha_{1}(\mathtt{i}|_{n})}\\ &\leqslant\lim_{n\to\infty}\frac{\log\mu([\mathtt{i}|_{n}])}{\log\alpha_{1}(\mathtt{i}|_{n})}=-\frac{h(\mu,\nu)}{\lambda_{1}(\nu)}=\tau^{\prime}(q)\end{split}$ (6.6)
for $\nu$-almost all $\mathtt{i}\in\Sigma$. By (6.5), Proposition 6.1, and (6.6), we thus get
$q\tau^{\prime}(q)-\tau(q)=\operatorname{\underline{dim}_{H}}(\pi_{*}\nu)\leqslant\operatorname{dim_{H}}(X_{t})\leqslant qt-\tau(q)\leqslant q\tau^{\prime}(q)-\tau(q)$
as claimed. ∎
Proposition 6.4.
Let $X$ be dominated irreducible planar self-affine set satisfying the strong separation condition, $\mu\in\mathcal{M}(\Sigma)$ be a fully supported quasi-Bernoulli measure, and let $q>0$ be such that $q\neq 1$. If $\tau(q)/(q-1)$ and $q\tau^{\prime}(q)-\tau(q)$ are contained in $(0,1)$, and there exist constants $\varrho_{3}>0$ and $C\geqslant 1$ such that
@@
The proof runs almost verbatim to the proof of Proposition 6.3: In (6.6), the assumption (6.7) guarantees an equality and hence, $\pi_{*}\nu(X_{\tau^{\prime}(q)})=1$ and $\operatorname{dim_{H}}(X_{\tau^{\prime}(q)})\geqslant\operatorname{\underline{dim}_{H}}(\pi_{*}\nu)$, where $\nu\in\mathcal{M}_{\sigma}(\Sigma)$ is the equilibrium state for $\psi^{q,\tau(q)/(q-1)}$. As (6.5) gives $\operatorname{\underline{dim}_{H}}(\pi_{*}\nu)=q\tau^{\prime}(q)-\tau(q)$, the proof is finished. ∎
@@
Let $X$ be dominated irreducible planar self-affine set satisfying the strong separation condition, $\mu\in\mathcal{M}(\Sigma)$ be a fully supported quasi-Bernoulli measure, and $q>0$ is such that $q\neq 1$. If $\tau(q)/(q-1)$ and $q\tau^{\prime}(q)-\tau(q)$ are contained in $(1,2)$, and
@@
We again follow the proof of Proposition 6.3. Since $\nu$ is quasi-Bernoulli, (5.1) and Theorem 4.2 imply also in this case that
@@
The assumption (6.8), Theorem 5.2, and Proposition 4.1 imply that
$\operatorname{dim_{loc}}(\pi_{*}\mu,\pi(\mathtt{i}))=\min\{2,\operatorname{dim_{L}}(\mu,\nu)\}=\min\{2,\tau^{\prime}(q)\}$
for $\nu$-almost all $\mathtt{i}\in\Sigma$. Writing $t=\min\{2,\tau^{\prime}(q)\}$, we thus have $\pi_{*}\nu(X_{t})=1$ and $\operatorname{dim_{H}}(X_{\tau^{\prime}(q)})\geqslant\operatorname{\underline{dim}_{H}}(\pi_{*}\nu)$ as claimed. ∎
@@
whenever $10<q<$ 1. The projective strong separation condition (1.2) and the compactness of both the self-affine set $X$ and the set of Furstenberg directions $X_{F}$ guarantee that
@@
By domination, it follows from [7, Lemma 2.8] that there exists a constant $D\geqslant 1$ such that
@@
whenever $10<q<$ 1. Since for each $V\in X_{F}$ there exists a Lebesgue integrable function $g_{V}\colon V^{\bot}\to\mathbb{R}$ such that $C=\sup_{V\in X_{F}}\|g_{V}\|_{\infty}<\infty$ and
@@ Acknowledgement.
In retrospect, AK acknowledges the discussions with Thomas Jordan, Pablo Shmerkin, and Károly Simon from around 2012, which provided a motivating catalyst and shaped the trajectory of the present work.

### v2

@@ 6 Multifractal Formalism
In this section, we extend the symbolic multifractal formalism results of Section 4 to obtain results in terms of the Hausdorff dimension of the sets
@@
This section contains several short propositions, some of which are quite similar; the proofs of Theorems 1.3 and 1.4 are presented after them.
@@
Let $X$ be a dominated planar self-affine set satisfying the strong separation condition, $\mu\in\mathcal{M}(\Sigma)$ be a fully supported quasi-Bernoulli measure, $0<q<1$, and $s\geqslant 0$. Then
@@
In general, the result follows from [18, Theorem 6.2(a)] and [31, Proposition 2.5(iv)]. Our version follows directly from [10, Lemmas 5.1 and 5.3] and the comments after the lemmas. The restriction $0<q<1$ arises because, in the language of [10, Lemma 5.1], the inequality between $\tau(\mu,q)$ and $\tau(q)$ changes sign at $q=1$. ∎
The arguments of the following proposition are an adjustment of those in [26, Theorem 4.1].
@@
$\{B(\pi(\mathtt{i}),\varrho\alpha_{m}(\mathtt{i}|_{n})):\;$ $\mathtt{i}\in\Sigma\text{ and }n\geqslant n_{\delta,\kappa}\text{ are such that}$
@@
is a Vitali covering of $X_{s}$. By the Vitali covering theorem (see e.g. [15, Theorem 1.10]), the above family contains a countable sub-family $\mathcal{V}$ of pairwise disjoint balls satisfying
@@
Since the elements of $\mathcal{V}$ are pairwise disjoint, a simple volume argument shows that, under the assumption 1, $\mathcal{V}_{\mathtt{j}}\leqslant c$, and under the assumption 2, $\mathcal{V}_{\mathtt{j}}\leqslant c\frac{\alpha_{1}(\mathtt{j})}{\alpha_{2}(\mathtt{j})}$ for some constant $c>0$ not depending on $\mathtt{j}$. Notice that, as $q>0$, the definition of $\mathcal{V}$ yields
@@
By (6.3), the assumption 1, the fact that $\mathcal{V}_{\mathtt{j}}\leqslant c$, and (6.1), we have
@@
for all $\mathtt{i}\in\Sigma$, $V\in X_{F}$, and $n\in\mathbb{N}$. Therefore, by (6.3), (6.4), the fact that $\mathcal{V}_{\mathtt{j}}\leqslant c\frac{\alpha_{1}(\mathtt{j})}{\alpha_{2}(\mathtt{j})}$, and (6.1), we similarly get
@@
Let $X$ be a dominated planar self-affine set satisfying the strong separation condition such that $X_{F}$ is not a singleton, $\mu\in\mathcal{M}(\Sigma)$ be a fully supported quasi-Bernoulli measure, and let $q>0$ be such that $q\neq 1$. If $\tau(q)/(q-1)$ and $q\tau^{\prime}(q)-\tau(q)$ are contained in $(0,1)$, and there exist constants $\varrho_{3}>0$ and $C\geqslant 1$ such that
@@
Let $\nu$ be the equilibrium state for $\psi^{q,\tau(q)/(q-1)}$. The containment $\varphi_{\mathtt{i}|_{n}}(X)\subset B(\pi(\mathtt{i}),\operatorname{diam}(X)\alpha_{1}(\mathtt{i}|_{n}))$ and (6.5) imply
$\frac{1}{C}\pi_{*}\mu(B(\pi(\mathtt{i}),\varrho_{3}\alpha_{1}(\mathtt{i}|_{n})))\leqslant\mu([\mathtt{i}|_{n}])\leqslant\pi_{*}\mu(B(\pi(\mathtt{i}),\operatorname{diam}(X)\alpha_{1}(\mathtt{i}|_{n}))).$
Since $\log C/\log\alpha_{1}(\mathtt{i}|_{n})\to 0$ for $\nu$-almost all $\mathtt{i}$, Lemma 2.1 and Proposition 4.1 yield $\operatorname{dim_{loc}}(\pi_{*}\mu,\pi(\mathtt{i}))=\tau^{\prime}(q)$ for $\nu$-almost all $\mathtt{i}$. Hence $\pi_{*}\nu(X_{\tau^{\prime}(q)})=1$. Since $\nu$ is quasi-Bernoulli, (5.2) and Theorem 4.2 imply $\operatorname{\underline{dim}_{H}}(\pi_{*}\nu)=q\tau^{\prime}(q)-\tau(q)$. Therefore
$\operatorname{dim_{H}}(X_{\tau^{\prime}(q)})\geqslant\operatorname{\underline{dim}_{H}}(\pi_{*}\nu)=q\tau^{\prime}(q)-\tau(q)$
and the proof is finished. ∎
@@
Let $X$ be a dominated planar self-affine set satisfying the strong separation condition such that $X_{F}$ is not a singleton, $\mu\in\mathcal{M}(\Sigma)$ be a fully supported quasi-Bernoulli measure, and $q>0$ is such that $q\neq 1$. If $\tau(q)/(q-1)$ and $q\tau^{\prime}(q)-\tau(q)$ are contained in $(1,2)$, and
@@
Set $s=\tau(q)/(q-1)$ and $t=q\tau^{\prime}(q)-\tau(q)$. By assumption, $s,t\in(1,2)$ and therefore
$\tau^{\prime}(q)=\frac{t+(q-1)s}{q}\in(1,2).$
Since $\nu$ is quasi-Bernoulli, (5.2) and Theorem 4.2 imply that
@@
Proposition 4.1 gives $\operatorname{dim_{L}}(\mu,\nu)=\tau^{\prime}(q)$. Therefore, the assumption (6.6) and Theorem 5.2 imply that
$\operatorname{dim_{loc}}(\pi_{*}\mu,\pi(\mathtt{i}))=\min\{2,\operatorname{dim_{L}}(\mu,\nu)\}=\tau^{\prime}(q)$
for $\nu$-almost all $\mathtt{i}\in\Sigma$. Thus $\pi_{*}\nu(X_{\tau^{\prime}(q)})=1$ and so $\operatorname{dim_{H}}(X_{\tau^{\prime}(q)})\geqslant\operatorname{\underline{dim}_{H}}(\pi_{*}\nu)$ as claimed. ∎
@@
whenever $0<q<1$. The projective strong separation condition (1.2) and the compactness of both the self-affine set $X$ and the set of Furstenberg directions $X_{F}$ guarantee that
@@
By domination, it follows from [5, Lemma 2.8] that there exists a constant $D\geqslant 1$ such that
@@
whenever $0<q<1$. Since for each $V\in X_{F}$ there exists a Lebesgue integrable function $g_{V}\colon V^{\bot}\to\mathbb{R}$ such that $C=\sup_{V\in X_{F}}\|g_{V}\|_{\infty}<\infty$ and
@@
To finish the discussion, let us prove Proposition 1.5. The Fourier transform of a finite Borel measure $\mu$ on $\mathbb{R}^{2}$ is defined by
$\hat{\mu}(\xi)=\int_{\mathbb{R}^{2}}e^{i\langle\xi,x\rangle}\,\mathrm{d}\mu(x)$
for all $\xi\in\mathbb{R}^{2}$ and the Fourier transform of an integrable function $\psi\colon\mathbb{R}^{d}\to\mathbb{R}$ by
$\hat{\psi}(\xi)=\int_{\mathbb{R}^{d}}\psi(\mathbf{v})e^{i\langle\xi,\mathbf{v}\rangle}\,\mathrm{d}\mathbf{v}$
for all $\xi\in\mathbb{R}^{d}$, where $\langle\,\cdot\,$$\cdot\,\rangle$ denotes the standard inner product on $\mathbb{R}^{d}$. If $\psi$ is a smooth function with compact support on $\mathbb{R}^{d}$, then it follows from the Riemann-Lebesgue lemma and integration by parts that for every $M$ there is a constant $C_{M}>0$ such that
$|\hat{\psi}(\xi)|\leqslant\frac{C_{M}}{(1+|\xi|)^{M}}$ (6.10)
for all $\xi\in\mathbb{R}^{d}$. There are intimate connections between the regularity of a measure and the decay rate of its Fourier transform. In particular, if $\mu$ is a finite Borel measure with compact support on $\mathbb{R}^{2}$ such that
$\int_{\mathbb{R}^{2}}|\xi|^{\beta}|\hat{\mu}(\xi)|^{2}\,\mathrm{d}\xi<\infty$ (6.11)
for some $\beta>2$, then, by [27, Theorem 5.4], $\mu$ is absolutely continuous with bounded continuous density.
Proof of Proposition 1.5.
Assuming that $\mu$ is a fully supported self-affine measure satisfying
$\sum_{i=1}^{N}|c_{i}|^{-s}\mu(\varphi_{i}(X))^{2}<1$ (6.12)
for some $s>3$, the proof requires only minor adjustments to that of [8, Proposition 4.2]: we replace the equilibrium state with $\mu$ and invoke assumption (6.12) at the conclusion to ensure the finiteness of the integral. The key steps are outlined below. We indicate at the end of the proof how the extension to quasi-Bernoulli measures is achieved.
We write $\mu_{\mathbf{v}}$ to emphasize the dependence of the self-affine measure on $\mathbf{v}=(v_{1}^{2},\ldots,v_{N}^{2})$. Let $s>3$ be such that (6.12) holds and choose $2<\beta<s-1$. Choose $h\colon\mathbb{RP}^{1}\to[0,\infty)$ to be a compactly supported continuous density function uniformly separated away from zero on the smallest projective interval [C,C][-C,C] containing $X_{F}$. We identify $\mathbb{RP}^{1}$ with $[0,1)$. Recalling (6.10), let $C_{1}>0$ be such that
$|\hat{h}(\xi)|\leqslant\frac{C_{1}}{(1+|\xi|)^{s}}$
for all $\xi\in\mathbb{R}$. We define a compactly supported measure $\nu_{\mathbf{v}}$ on $\mathbb{R}^{2}$ by
$\nu_{\mathbf{v}}(A)=\iint_{A}h(x)\,\mathrm{d}(\operatorname{proj}_{x})_{*}\mu_{\mathbf{v}}(y)\,\mathrm{d}x$
for all Borel sets $A\subset\mathbb{R}^{2}$, where $\operatorname{proj}_{x}$ with $\mathrm{im}\operatorname{proj}_{x}=\mathbb{R}$ corresponds to the orthogonal projection $\operatorname{proj}_{V^{\perp}}$ via the identification. Let $\psi\colon\mathbb{R}^{N}\to[0,\infty)$ be a compactly supported density function and, by (6.10), let $C_{2}>0$ be such that
$|\hat{\psi}(\xi)|\leqslant\frac{C_{2}}{(1+\|\xi\|)^{s}}$
for all $\xi\in\mathbb{R}^{N}$. By the proof of [8, Proposition 4.2] and (6.12), there is a constant $C_{3}>0$ such that
$\begin{split}\biggl|\int_{\mathbb{R}^{N}}&\int_{\mathbb{R}^{2}}|\xi|^{\beta}|\hat{\nu}_{\mathbf{v}}(\xi)|^{2}\,\mathrm{d}\xi\;\psi(\mathbf{v})\,\mathrm{d}\mathbf{v}\biggr|\leqslant\cdots\\ &\leqslant C_{3}\int_{\mathbb{R}^{2}}\frac{|(\xi_{1},\xi_{2})|^{\beta}}{(1+|\xi_{1}|)^{s}(1+|\xi_{2}|)^{s}}\,\mathrm{d}\xi_{1}\,\mathrm{d}\xi_{2}\sum_{n\in\mathbb{N}}\biggl(\sum_{i=1}^{N}\frac{\mu(\varphi_{i}(X))^{2}}{|c_{i}|^{s}}\biggr)^{n}\\ &<\infty,\end{split}$ (6.13)
where the calculation omitted by the ellipsis is identical to that in the proof of [8, Proposition 4.2]. Therefore, for Lebesgue almost every choice of $\mathbf{v}$, by (6.11), $\nu_{\mathbf{v}}$ is absolutely continuous with bounded continuous density, i.e., there exists a continuous $g_{\mathbf{v}}\colon\mathbb{R}^{2}\to[0,\infty)$ such that $\nu_{\mathbf{v}}(A)=\int_{A}g_{\mathbf{v}}(x,y)\,\mathrm{d}(x,y)$ for all Borel sets $A\subset\mathbb{R}^{2}$. Since $h$ is uniformly separated away from zero, it follows that $(\operatorname{proj}_{x})_{*}\mu$ is absolutely continuous with uniformly bounded continuous density $g_{\mathbf{v}}(x,y)/h(x)$.
The proof extends to quasi-Bernoulli measures $\mu$ by invoking the assumption (1.3) at the calculation (6.13) to ensure the finiteness of the integral in question. ∎
@@ Acknowledgement.
In retrospect, AK gratefully acknowledges discussions with Thomas Jordan, Pablo Shmerkin, and Károly Simon around 2012, which served as a motivating catalyst and significantly shaped the trajectory of this work. Additionally, he expresses gratitude to Balázs Bárány for valuable conversations related to Proposition 1.5.
Declaration. The authors have no relevant financial or non-financial interests to disclose. This work was not supported by grants from any funding bodies.